OpenStudy (anonymous):

Toni just started a new job earning 24.85/hour. Find and interpret the percentile for toni's earnings

OpenStudy (anonymous):

mean= 17.50 standard deviation= 3.50

OpenStudy (anonymous):

First we need to find a z score. Z= (X-mean)/(Std Dev) = (24.85-17.5)/3.50= 2.1 Then we go to our table for a standard normal distribution and check the p value. We get 0.9821. This means Toni is earning more than 98.21% of other people.